# Minimal criteria for continuous-variable genuine multipartite entanglement

### v1

@@ Abstract
pacs:
03.67.-a
I Introduction
At the very foundations of the quantum information theory lies the effort to explain possible relationships between a pair of quantum systems. These relationships manifest itself through various forms of correlations which we try to capture by intuitive models complying better with our idea of how the nature ought to work. Then we can classify states of a composite system into a nested family of sets according to whether the state admits a separable model [1], local hidden state model [2] or local hidden variable model [3]. The absence of the model then points at a closer relationship between the subsystems, which is referred to as entanglement, steering and non-locality, and which can be used as a resource for quantum communication and computing [4].
The complexity of the relationships is significantly increased in systems consisting of three or more subsystems, called commonly as the multipartite systems. Here, a particularly important role is played by the genuine multipartite entanglement (GME) [5], which is a synonym for truly multipartite inseparability of the subsystems. This means, in other words, that a state carrying GME cannot be modelled as a statistical mixture of product states. Physically, preparation of a GME state requires interaction of all subsystems involved [6].
The significance of GME states follows from the fact that they stay behind violation of the statistical [7] and non-statistical [8] constraints imposed by local hidden variable models, complex behaviour of strongly correlated systems [9] and tests of multiparticle decoherence models [10]. Besides being of fundamental interest, the GME states also find applications in measurement-based quantum computing [11], quantum simulations [12], and quantum metrology [13].
@@ I Introduction
Attempts to prepare a large GME state are proceeding in two directions. One uses systems with a two-dimensional Hilbert state space (qubits), while the other relies on systems with an infinite-dimensional Hilbert state space, the so-called continuous-variable (CV) systems. Nowadays, preparation of the qubit GME states is limited to the medium-scale systems. This includes preparation of the 14 ion qubits in the Greenberger-Horne-Zeilinger (GHZ) state [10], 18-qubit photonic GHZ state [16] or 51-qubit cluster state on a superconducting quantum processor [17] to name at least few of them.
@@
There are two basic ways in which the CV GME can be detected. The first one utilizes the GME witness in the space of covariance matrices [26]. It requires knowledge of the entire covariance matrix, or its entire position and momentum part, and so far it has been used exclusively for verification of the GME in the theoretical states [27, 28]. The second approach is based on simpler CV GME criteria [19, 29, 22, 30], which are more transparent and more economical in terms of the number of needed measurements. All the criteria are built on the van Loock-Furusawa criterion of full inseparability [31], which generalizes Duan’s et. al. criterion of two-mode entanglement [32] to the multi-mode case. The considered criteria are usually designed to detect in a simple way some paradigmatic GME Gaussian states and therefore they are of the form of a single inequality for the second moments of the quadrature operators. As an example can serve us the sum version of the criterion [19], according to which a three-mode state is GME if the inequality
@@
is violated. The criterion exhibits a common feature of the overwhelming majority of the practically used GME criteria. Namely, some (more often all [29, 20]) occurring variances contain a global combination of all position or momentum quadratures. For the three-mode criterion (I) the combination in question is the total momentum quadrature $p_{1}+p_{2}+p_{3}$, whereas more general $N$-mode criteria typically contain the weighted combination $\sum_{i=1}^{N}g_{i}p_{i}$. Consequently, utilization of such the criterion requires knowledge of all independent elements of the momentum part of the covariance matrix, the number of which scales quadratically with $N$. Guided by the effort to detect GME of the largest possible state, we can ask ourselves whether this quadratic growth can be slowed down, e.g., just to the linear growth. Finding a positive answer to this question would undoubtedly be an important result because then the GME certification of some CV states would require a significantly smaller number of measurements. However, such a result would be rather counterintuitive because the respective criterion then could not contain any global combination mentioned above, which might seem to be indispensable for the detection of the global property of the GME.
Interestingly, this intuition is false. Namely, there are two similar criteria of GME, which nevertheless contain variances of combinations of only pairs of position and momentum quadratures. More precisely, one criterion is $N$-mode and it is given by [22, Eq. (25)],
@@
where $f(N)$ is a function of $N$, whereas the other is four-mode and employs variances of the same combinations (up to the sign) of quadrature operators [30, Eqs. (44) and (47)]. Importantly, both the criteria contain combinations for all possible pairs of the modes. Thus as before, the number of the elements of the covariance matrix one needs for application of both the latter criteria grows quadratically with the number of modes $N$.
@@
In this paper we derive such the simplified CV GME criteria for $N$ modes. Similar to the known criteria also the proposed criteria have the standard form akin to the product and sum uncertainty relation for elements of the covariance matrix and like the criteria [22, Eq. (25)] and [30, Eqs. (44) and (47)] they include at most two-mode quadrature combinations. However, unlike any of the existing CV GME criteria they contain the minimum number of the combinations. The number is minimal because it is equal to the number of elements in the minimal set of two-mode marginal covariance matrices [33, 35], which is needed for detection of GME. Since the number of independent covariance matrix elements contained in the minimal set is equal to $47N-$ 4 [35] and thus grows linearly with $N$, the number of elements of the covariance matrix one needs to know for application of the proposed criteria also scales only linearly with $N$. As every minimal set can be represented by a tree graph, each of the presented criterion corresponds to a tree. This allowed us to developed a direct method of construction of the criteria, which is based solely on the structure of the underlying tree. We also demonstrate practicality of our criteria by finding analytical as well as numerical examples of Gaussian states whose GME is detected by the criteria for almost all configurations of up to six modes.
The proposed criteria are most economic as far as the number of required measurements is concerned and thus they can serve for verification of GME in large multi-mode states. They are suitable for detection of the CV GME in instances when only elements of two-mode marginal covariance matrices belonging to some minimal set are know. Viewed from a different perspective, the derived criteria also represent a sufficient condition for a set of two-mode marginals to be compatible with a GME state.
@@
The paper is structured as follows. In Sec. II we give a brief introduction into the concept of GME, whereas Sec. III is dedicated to the entanglement in CV systems. Section IV contains derivation of the minimal criteria for CV GME and the Gaussian GME states which are detected by the criteria are constructed in Sec. V. Conclusions are in Sec. VI.
II Genuine multipartite entanglement
@@
We deal with $N$ quantum systems $1,2,\ldots,N$, which we collect into the set =$\mathcal{M}=\{1,2,\ldots,N\}$. Consider further the split of the set $\mathcal{M}$ into two disjoint nonempty subsets $I_{k}=\{i_{1},i_{2},\ldots,i_{l}\}$, $i_{n}\in\mathcal{M}$, $n=1,2,\ldots,l$, $0<l<N$, and $J_{k}=\mathcal{M}\setminus I_{k}$. Here, the index $k=1,2,\ldots,K_{N}$ labels all different inequivalent splits of the set $\mathcal{M}$ and $K_{N}=2^{N-1}-1$ is the number of the splits. The splits shall be henceforth called as bipartite splits and denoted as $I_{k}|J_{k}$. We then say that a state $\rho$ of the considered system is separable with respect to the split $I_{k}|J_{k}$ if it can be expressed as
@@
States of systems made up of only two subsystems ($N=2$) are either entangled or separable, but in systems consisting of more subsystems ($N>2$) different types of entanglement exist [36, 37]. Various forms of the multipartite entanglement can be illustrated on the system of just three subsystems $1,2$ and $3$. In this case, there are three bipartite splits 1|23,2|311|23,2|311 | 23 , 2 | 31 and $3|12$, as well as a tripartite split $1|2|3$, and one can classify all states into five separability classes [36, 37] depending on the number of the splits with respect to which a given state is separable. Clearly, from the point of view of this classification the strongest form of multipartite entanglement is carried by the so-called fully inseparable states, which are entangled across all three bipartite splits.
A specific feature of most important fully inseparable states is that their preparation requires interaction of all three subsystems. However, this is not a common property of all fully inseparable states. Namely, consider a class of states which can be prepared by letting to interact randomly only pairs of subsystems prepared in a triple product state. In this way, one can prepare only a convex mixture of states which are separable across some bipartite split, which is commonly denoted as a biseparable state. Interestingly, there are qubit biseparable states which are fully inseparable [38]. To distinguish truly multipartite entangled states requiring a global operation on the whole system from other multipartite states, the concept of GME has been introduced as follows. Generalizing the notion of a biseparable state to $N$-partite system,
@@
III Continuous-variable entanglement
We assume that the considered systems $1,2,\ldots,N$ are the CV systems. Each of the systems is then fully characterized by the position and momentum operators $x_{j}$ and $p_{j}$ ($[x_{j},p_{k}]=i\delta_{j,k}$), $j=1,2,\ldots,N$, respectively. From now on, we work with the optical platform, where the systems are realized by modes and the position and momentum operators by the respective quadrature operators. It is further convenient to introduce the column vector $\xi=\left(\xi_{x}^{T},\xi_{p}^{T}\right)^{T}$, where $\xi_{x}=\left(x_{1},x_{2},\dots,x_{N}\right)^{T}$ and $\xi_{p}=\left(p_{1},p_{2},\dots,p_{N}\right)^{T}$, which allows us to express the quadrature commutation rules in the compact form $[\xi_{j},\xi_{k}]=i(\Omega_{N})_{jk}$ with
[…]
@@ III Continuous-variable entanglement
$\rho^{\rm test}$ $=$ $\frac{1}{3}\left(\rho^{\rm TMSV}_{12}\otimes|0\rangle_{3}\langle 0|+\rho_{13}^{\rm TMSV}\otimes|0\rangle_{2}\langle 0|\right.$ (13)
$\left.+|0\rangle_{1}\langle 0|\otimes\rho_{23}^{\rm TMSV}\right),$
@@
[…]
one can show (see Appendix A for details of the calculation) that for $r\in(0,1.24)$ the state (13) is at the same time entangled across all three bipartite splits 1|23,2|311|23,2|311 | 23 , 2 | 31 and $3|12$, and thus it is fully inseparable as required.
@@
IV Minimal criteria for genuine multipartite entanglement
@@
IV.1 Left-hand side of the criterion
At the outset we derive the LHS of the sought GME criterion. Similar to other entanglement criteria we rule out combinations mixing position and momentum quadratures by assuming $Z=Z^{x}\oplus Z^{p}$, where $Z^{x}$ and $Z^{p}$ are real, symmetric, and positive-semidefinite $N\times N$ matrices. Inspired by the position part of the criterion (I), which involves combinations $x_{1}-x_{2}$ and $x_{2}-x_{3}$, we further seek the criteria containing the least number of only two-mode quadrature combinations. We therefore require the criteria to utilize only the minimal number of two-mode reduced CMs (marginal CMs), which are needed for detection of GME. The marginal CMs comprise the so-called minimal set of marginals and its structure is already known [33, 35]. The set has to contain all modes and one cannot divide it into a subset and its complement without having a common mode.
@@ IV.1 Left-hand side of the criterion
It remains to rewrite the RHS of Eq. (19) in terms of variances of two-mode quadrature combinations. Since the witness matrices $Z_{A}^{x,p}$ are always real, symmetric and positive-semidefinite, one can use the Cholesky decomposition and express them as $Z_{A}^{\alpha}=L_{A}^{\alpha}(L_{A}^{\alpha})^{T}$, where $L_{A}^{\alpha}$, $\alpha=x,p$, are real lower triangular $N\times N$ matrices with non-negative diagonal elements [49, 50, 51]. Like in the case of the matrix $Z_{A}^{\alpha}$ the structure of the Cholesky matrix $L_{A}^{\alpha}$ also depends on the respective adjacency matrix $A$, which is expressed by the lower index $A$. In what follows, the lower index is dropped for brevity, i.e., from now by the symbol $L^{\alpha}$ we understand a Cholesky matrix corresponding some adjacency matrix $A$.
@@
$\mathrm{Tr}[\gamma^{\alpha}Z_{A}^{\alpha}]$ $=$ $\mathrm{Tr}[\gamma^{\alpha}L^{\alpha}(L^{\alpha})^{T}]=2\sum_{i=1}^{N}\langle\{\Delta[(L^{\alpha})^{T}\xi_{\alpha}]_{i}\}^{2}\rangle$ (20)
$=$ $2\sum_{i=1}^{N}\langle\left(\Delta u_{i}^{\alpha}\right)^{2}\rangle\equiv 2U^{\alpha},$
@@
At the outset we show how the requirement that quadrature combinations (21) contain at most two terms manifests itself on the level of the respective adjacency matrix $A$. Recall first, that we seek a necessary condition for biseparability of the form $f_{\gamma}(L^{x},L^{p})\geq g(L^{x},L^{p})$. On the LHS there is a function $f_{\gamma}(L^{x},L^{p})$ which depends on CM $\gamma$ of the tested state and Cholesky matrices $L^{x,p}$ of some witness matrices $Z_{A}^{x,p}$. The RHS of the condition, which we derive in the next section, imposes a lower bound on the LHS for all biseparable states, and it is formed by a function $g(L^{x},L^{p})$ depending only on the matrices $L^{x,p}$. The biseparability condition then can be used for testing of the presence of GME in a given CM $\gamma$ by numerical minimization of the difference $f_{\gamma}(L^{x},L^{p})-g(L^{x},L^{p})$ over the matrices $L^{x,p}$. If the minimum is negative, i.e., the biseparability condition is violated, then the state with CM $\gamma$ is inevitably GME. Here, we are interested only in a generic form of the biseparability condition capable of detecting the largest number of GME states from a minimal set of marginal CMs. Clearly, such a condition is obtained if the function $f_{\gamma}(L^{x},L^{p})$ depends on the maximum number of independent variables, i.e., if the Cholesky matrices $L^{x,p}$ corresponding to the minimal set have the maximum number of non-zero elements. Working with the positive-semidefinite matrices $Z_{A}^{x,p}$, we would have to take into account that some diagonal elements of the Cholesky matrices $L^{x,p}$ can be equal to zero [51, p. 8.3] thereby reducing the number of the variables over which we will optimize. For this reason we assume from now that the matrices $Z_{A}^{x,p}$ are positive-definite which implies that the corresponding Cholesky matrices $L^{x,p}$ possess strictly positive diagonal elements which can be expressed as [50, Theorem 4.2.5]
@@
Now, let us rewrite the first $N-1$ combinations (21) in the compact form
@@
[…]
@@
are $(N-i)\times 1$ column vectors comprising lower triangular part without the main diagonal of the matrix $L^{\alpha}$. Hence, we see that the operators (IV.1) will contain at most two non-zero terms if each of the vectors $L^{\alpha}_{i+1:N,i}$ will have at most one non-zero component. For a positive-definite matrix $Z_{A}^{\alpha}$ the vectors can be further expressed as [50, Theorem 4.2.5]
$L^{\alpha}_{2:N,1}=\frac{(Z_{A}^{\alpha})_{2:N,1}}{L_{11}^{\alpha}}$ (26)
and
$L^{\alpha}_{i+1:N,i}=\frac{1}{L_{ii}^{\alpha}}\left[(Z_{A}^{\alpha})_{i+1:N,i}-\sum_{j=1}^{i-1}L_{ij}^{\alpha}L^{\alpha}_{i+1:N,j}\right]$
for $i=2,3,\ldots,N-1$. From Eq. (26) it then follows that the vector $L^{\alpha}_{2:N,1}$ will contain at most one non-zero component if the vector $(Z_{A}^{\alpha})_{2:N,1}$ has at most one non-zero component. Moving to the vector $L^{\alpha}_{3:N,2}$, Eq. (IV.1) with $i=2$, we see that if the vector $(Z_{A}^{\alpha})_{2:N,1}$ has at most one non-zero component, then the second term in the square brackets vanishes because it contains products of different components of the vector $(Z_{A}^{\alpha})_{2:N,1}$, and $L^{\alpha}_{3:N,2}=(Z_{A}^{\alpha})_{3:N,2}/L^{\alpha}_{22}$ according to Eq. (IV.1). Again, in order the vector $L^{\alpha}_{3:N,2}$ to have at most one non-zero component, the vector $(Z_{A}^{\alpha})_{3:N,2}$ has to possess at most one non-zero component. Repeating the same procedure $(i-1)$-times we see that the vectors $L^{\alpha}_{j+1:N,j}$, $j=1,2,\ldots,i-1$ will have at most one non-zero component if the same holds for vectors $(Z_{A}^{\alpha})_{j+1:N,j}$, $j=1,2,\ldots,i-1$. If this is the case, then the sum $\sum_{j=1}^{i-1}L_{ij}^{\alpha}L^{\alpha}_{i+1:N,j}$ in Eq. (IV.1) vanishes, because each term $L_{ij}^{\alpha}L^{\alpha}_{i+1:N,j}$ is a vector with components given by a product of different components of the vector $(Z_{A}^{\alpha})_{j+1:N,j}$. The $i$-th vector (IV.1) of the Cholesky matrix $L^{\alpha}$ then reduces to
$L^{\alpha}_{i+1:N,i}=\frac{(Z_{A}^{\alpha})_{i+1:N,i}}{L_{ii}^{\alpha}}$ (28)
and again it has at most one non-zero element if the vector $(Z_{A}^{\alpha})_{i+1:N,i}$ contains at most one non-zero element.
Summarizing the obtained results we see that if we require all the quadrature combinations (21) to contain at most two non-zero terms, then all vectors $(Z_{A}^{\alpha})_{i+1:N,i}$, $i=1,2,\ldots,N-1$ have to possess at most one non-zero element. Making further use of the formula (18) we can express the latter vectors in the form

### v3

@@ Abstract
1 Introduction
At the very foundations of the quantum information theory lies the effort to explain possible relationships between a pair of quantum systems. These relationships manifest itself through various forms of correlations which we try to capture by intuitive models complying better with our idea of how the nature ought to work. Then we can classify states of a composite system into a nested family of sets according to whether the state admits a separable model [1, 2], local hidden state model [3, 4] or local hidden variable model [5, 6]. The absence of the model then points at a closer relationship between the subsystems, which is referred to as entanglement, steering and non-locality, and which can be used as a resource for quantum communication and computing [7].
The complexity of the relationships is significantly increased in systems consisting of three or more subsystems, called commonly as the multipartite systems. From the point of view of quantum entanglement, states of multipartite systems can be classified into two disjoint sets. One is the set of so-called biseparable states, which can be prepared by statistical mixing of states that are separable with respect to (generally different) partitions of all subsystems into two groups. The second is then the set of states that cannot be prepared in this way, which we refer to as genuine multipartite entangled (GME) [8].
Genuine multiparticle entanglement is of great importance both in terms of fundamental research and applications. The GME carried by the Greenberger-Horne-Zeilinger (GHZ) state exhibits a strong contradiction with locally realistic theories, which can be captured without the use of the Bell-type statistical inequalities [9, 10]. The same state also serves as an ideal testbed for the analysis of the scaling law of the coherence decay of a multiparticle quantum register [11]. Early applications of GME included quantum secret sharing protocol [12] or assisted quantum teleportation schemes [13, 14]. In newer applications, GME is used to increase the precision of phase estimation in interferometry [15, 16] and it also plays a key role in the measurement-based model of quantum computing [17, 18], in which the computation is performed using a sequence of local single-site measurements on a large GME state. Last but not least, GME is also a resource for multiparty quantum key distribution [19, 20].
@@ 1 Introduction
Attempts to prepare a large GME state are proceeding in two directions. One uses systems with a two-dimensional Hilbert state space (qubits), while the other relies on systems with an infinite-dimensional Hilbert state space, the so-called continuous-variable (CV) systems. Nowadays, preparation of the qubit GME states is limited to the medium-scale systems. This includes preparation of the 14 ion qubits in the GHZ state [11], 18-qubit photonic GHZ state [23] or 51-qubit cluster state on a superconducting quantum processor [24] to name at least few of them.
@@
There are two basic ways in which the CV GME can be detected. The first one utilizes the GME witness in the space of covariance matrices [33]. It requires knowledge of the entire covariance matrix, or its entire position and momentum part, and so far it has been used exclusively for verification of the GME in the theoretical states [34, 35]. The second approach is based on simpler CV GME criteria [26, 36, 29, 37], which are more transparent and more economical in terms of the number of needed measurements. All the criteria are built on the van Loock-Furusawa criterion of full inseparability [38], which generalizes Duan’s et. al. criterion of two-mode entanglement [39] and its bipartite generalization [40] to the multi-mode case. The considered criteria are usually designed to detect in a simple way some paradigmatic GME Gaussian states and therefore they are of the form of a single inequality for the second moments of the quadrature operators. As an example can serve us the sum version of the criterion [26], according to which a three-mode state is GME if the inequality
@@
is violated. The criterion exhibits a common feature of the overwhelming majority of the practically used GME criteria. Namely, some (more often all [36, 27]) occurring variances contain a global combination of all position or momentum quadratures. For the three-mode criterion (1) the combination in question is the total momentum quadrature $p_{1}+p_{2}+p_{3}$, whereas more general $N$-mode criteria typically contain the weighted combination $\sum_{i=1}^{N}g_{i}p_{i}$. Consequently, utilization of such the criterion requires knowledge of all independent elements of the momentum part of the covariance matrix, the number of which scales quadratically with $N$. Guided by the effort to detect GME of the largest possible state, we can ask ourselves whether this quadratic growth can be slowed down, e.g., just to the linear growth. Finding a positive answer to this question would undoubtedly be an important result because then the GME certification of some CV states would require a significantly smaller number of measurements. However, such a result would be rather counterintuitive because the respective criterion then could not contain any global combination mentioned above, which might seem to be indispensable for the detection of the global property of the GME.
Interestingly, this intuition is false. Namely, there are two similar criteria of GME, which nevertheless contain variances of combinations of only pairs of position and momentum quadratures. More precisely, one criterion is $N$-mode and it is given by [29, Eq. (25)],
@@
where $f(N)$ is a function of $N$, whereas the other is four-mode and employs variances of the same combinations (up to the sign) of quadrature operators [37, Eqs. (44) and (47)]. Importantly, both the criteria contain combinations for all possible pairs of the modes. Thus as before, the number of the elements of the covariance matrix one needs for application of both the latter criteria grows quadratically with the number of modes $N$.
@@
In this paper we derive such the simplified CV GME criteria for $N$ modes. Similar to the known criteria also the proposed criteria have the standard form akin to the product and sum uncertainty relation for elements of the covariance matrix and like the criteria [29, Eq. (25)] and [37, Eqs. (44) and (47)] they include at most two-mode quadrature combinations. However, unlike any of the existing CV GME criteria they contain the minimum number of the combinations. The number is minimal because it is equal to the number of elements in the minimal set of two-mode marginal covariance matrices [41, 43], which is needed for detection of GME. Since the number of independent covariance matrix elements contained in the minimal set is equal to 7N47N-4 [43] and thus grows linearly with $N$, the number of elements of the covariance matrix one needs to know for application of the proposed criteria also scales only linearly with $N$. As every minimal set can be represented by a tree graph, each of the presented criterion corresponds to a tree. This allowed us to develop a direct method of construction of the criteria, which is based solely on the structure of the underlying tree. We also demonstrate practicality of our criteria by finding analytical as well as numerical examples of Gaussian states whose GME is detected by the criteria for almost all configurations of up to six modes. Interestingly, our criterion detects the GME of a three-mode CV GHZ-like state from its two two-mode reduced density matrices. This is in sharp contrast to the qubit GHZ state for which detection of the GME from two-qubit marginals is impossible [44] and illustrates another difference between discrete and continuous quantum variables.
The proposed criteria are most economic as far as the number of required measurements is concerned and thus they can serve for verification of GME in large multi-mode states. They can be used to detect CV GME even in cases where only the elements of two-mode marginal covariance matrices belonging to some minimal set are available. From a different point of view, the derived criteria can be interpreted in the context of the so-called quantum marginal problem [45, 46]. In its broader sense, this problem addresses the question of determining the presence of some global property based on the knowledge of only a set of reduced density matrices [47, 30]. The task of finding the conditions under which the given reductions are compatible with the property under consideration is then known as the quantum marginal problem. From this perspective, each of the presented criteria can be viewed as a necessary condition for the corresponding minimal set of two-mode covariance matrices having a common global state to be compatible with a biseparable state.
@@
The paper is structured as follows. In Sec. 2 we give a brief introduction into the concept of GME, whereas Sec. 3 is dedicated to the entanglement in CV systems. Section 4 contains derivation of the minimal criteria for CV GME and the Gaussian GME states which are detected by the criteria are constructed in Sec. 5. Conclusions are in Sec. 6.
2 Genuine multipartite entanglement
@@
We deal with $N$ quantum systems $1,2,\ldots,N$, which we collect into the set $\mathcal{M}=\{1,2,\ldots,N\}$. Consider further the split of the set $\mathcal{M}$ into two disjoint nonempty subsets $I_{k}=\{i_{1},i_{2},\ldots,i_{l}\}$, $i_{n}\in\mathcal{M}$, $n=1,2,\ldots,l$, $0<l<N$, and $J_{k}=\mathcal{M}\setminus I_{k}$. Here, the index $k=1,2,\ldots,K_{N}$ labels all different inequivalent splits of the set $\mathcal{M}$ and $K_{N}=2^{N-1}-1$ is the number of the splits. The splits shall be henceforth called as bipartite splits and denoted as $I_{k}|J_{k}$. We then say that a state $\rho$ of the considered system is separable with respect to the split $I_{k}|J_{k}$ if it can be expressed as [1, 2]
@@
States of systems made up of only two subsystems ($N=2$) are either entangled or separable, but in systems consisting of more subsystems ($N>2$) different types of entanglement exist [48, 49]. Various forms of the multipartite entanglement can be illustrated on the system of just three subsystems $1,2$ and $3$. In this case, there are three bipartite splits $1|23,2|31$ and $3|12$, as well as a tripartite split $1|2|3$, and one can classify all states into five separability classes [48, 49] depending on whether they are separable with respect to tripartite split, or only with respect to none, one, two, or three bipartite splits. Clearly, from the point of view of this classification the strongest form of multipartite entanglement is carried by the so-called fully inseparable states, which are entangled across all three bipartite splits.
A specific feature of the most important fully inseparable states, such as the GHZ state, is that global quantum operation on all three subsystems is needed to prepare them. However, this is not a common property of all fully inseparable states. Namely, there is a class of fully inseparable states that can be prepared more easily by randomly mixing states that do not require such a global operation for their preparation [50]. To show this, consider a state $|\chi\rangle_{ijk}\equiv|\phi_{+}\rangle_{ij}|0\rangle_{k}$ of three qubits $i,j,k$, where $|\phi_{+}\rangle_{ij}=(|00\rangle_{ij}+|11\rangle_{ij})/\sqrt{2}$ is the Bell state, whose preparation obviously does not need a global three-qubit quantum operation. Let us now imagine that we randomly prepare with the same probability $1/3$ one of the states $|\chi\rangle_{123},|\chi\rangle_{132}$ and $|\chi\rangle_{231}$. Interestingly, the result is also a fully inseparable state [51]. To distinguish states that can be prepared by convex mixing of the product of states with respect to different bipartite splits, called biseparable states [8], from states that cannot be prepared in this way, the concept of GME has been introduced as follows. Generalizing the notion of a biseparable state to $N$-partite system,
@@
3 Continuous-variable entanglement
We assume that the considered systems $1,2,\ldots,N$ are the CV systems. Each of the systems is then fully characterized by the position and momentum operators $x_{j}$ and $p_{j}$ ($x_{j},p_{k}]=i\delta_{j,k}$), $j=1,2,\ldots,N$, respectively. From now on, we work with the optical platform, where the systems are realized by modes and the position and momentum operators by the respective quadrature operators. It is further convenient to introduce the column vector $\xi=\left(\xi_{x}^{T},\xi_{p}^{T}\right)^{T}$, where $\xi_{x}=\left(x_{1},x_{2},\dots,x_{N}\right)^{T}$ and $\xi_{p}=\left(p_{1},p_{2},\dots,p_{N}\right)^{T}$, which allows us to express the quadrature commutation rules in the compact form $\xi_{j},\xi_{k}]=i(\Omega_{N})_{jk}$ with
$\Omega_{N}=\begin{pmatrix}\mathbb{O}_{N}&\mathbbm{1}_{N}\\ -\mathbbm{1}_{N}&\mathbb{O}_{N}\end{pmatrix}$, (5)
where $\mathbbm{1}_{N}$ and $\mathbb{O}_{N}$ is the $N\times N$ identity matrix and zero matrix, respectively. A particularly important class of states of CV systems is given by the so-called Gaussian states, which are defined as states with a Gaussian-shaped Wigner phase-space distribution. For $N$ modes the states are therefore fully described by the $2N\times 1$ vector $\langle\xi\rangle=\mathrm{Tr}(\xi\rho)$ of the first moments and by the $2N\times 2N$ real symmetric covariance matrix (CM) $\gamma$ with entries $(\gamma)_{jk}=\langle\{\Delta\xi_{j},\Delta\xi_{k}\}\rangle$, where $\{A,B\}\equiv AB+BA$ is the anticommutator and $\Delta A\equiv A-\langle A\rangle$. From the definition of the CM it further follows that it obeys the uncertainty principle [54]
@@ 3 Continuous-variable entanglement
$\rho^{\rm test}=$ $\frac{1}{3}\left(\rho^{\rm TMSV}_{12}\otimes|0\rangle_{3}\langle 0|+\rho_{13}^{\rm TMSV}\otimes|0\rangle_{2}\langle 0|\right$.
$\left.+|0\rangle_{1}\langle 0|\otimes\rho_{23}^{\rm TMSV}\right),$ (11)
@@
$\gamma^{\rm test}=\frac{1}{3}(\gamma^{\rm TMSV}_{12}\oplus\mathbbm{1}_{3}+\gamma_{13}^{\rm TMSV}\oplus\mathbbm{1}_{2}+\mathbbm{1}_{1}\oplus\gamma_{23}^{\rm TMSV}),$ (12)
one can show (see Appendix A for details of the calculation) that for $r\in(0,1.24)$ the state (3) is at the same time entangled across all three bipartite splits $1|23,2|31$ and $3|12$, and thus it is fully inseparable as required.
@@
4 Minimal criteria for genuine multipartite entanglement
@@
4.1 Left-hand side of the criterion
At the outset we derive the LHS of the sought GME criterion. Similar to other entanglement criteria we rule out combinations mixing position and momentum quadratures by assuming $Z=Z^{x}\oplus Z^{p}$, where $Z^{x}$ and $Z^{p}$ are real, symmetric, and positive-semidefinite $N\times N$ matrices. Inspired by the position part of the criterion (1), which involves combinations $x_{1}-x_{2}$ and $x_{2}-x_{3}$, we further seek the criteria containing the least number of only two-mode quadrature combinations. We therefore require the criteria to utilize only the minimal number of two-mode reduced CMs (marginal CMs), which are needed for detection of GME. The marginal CMs comprise the so-called minimal set of marginals and its structure is already known [41, 43]. The set has to contain all modes and one cannot divide it into a subset and its complement without having a common mode.
@@ 4.1 Left-hand side of the criterion
It remains to rewrite the RHS of Eq. (17) in terms of variances of two-mode quadrature combinations. Since the witness matrices $Z_{A}^{x,p}$ are always real, symmetric and positive-semidefinite, one can use the Cholesky decomposition and express them as $Z_{A}^{\alpha}=L_{A}^{\alpha}(L_{A}^{\alpha})^{T}$, where $L_{A}^{\alpha}$, $\alpha=x,p$, are real lower triangular $N\times N$ matrices with non-negative diagonal elements [62, 63, 64]. Like in the case of the matrix $Z_{A}^{\alpha}$ the structure of the Cholesky matrix $L_{A}^{\alpha}$ also depends on the respective adjacency matrix $A$, which is expressed by the lower index $A$. In what follows, the lower index is dropped for brevity, i.e., from now by the symbol $L^{\alpha}$ we understand a Cholesky matrix corresponding to some adjacency matrix $A$.
@@
$\mathrm{Tr}[\gamma^{\alpha}Z_{A}^{\alpha}]=$ $\mathrm{Tr}[\gamma^{\alpha}L^{\alpha}(L^{\alpha})^{T}]$
$=$ $2\sum_{i=1}^{N}\langle\{\Delta[(L^{\alpha})^{T}\xi_{\alpha}]_{i}\}^{2}\rangle$
$=$ $2\sum_{i=1}^{N}\langle\left(\Delta u_{i}^{\alpha}\right)^{2}\rangle\equiv 2U^{\alpha},$ (18)
@@
At the outset we show how the requirement that quadrature combinations (19) contain at most two terms manifests itself on the level of the respective adjacency matrix $A$. Recall first, that we seek a necessary condition for biseparability of the form $f_{\gamma}(L^{x},L^{p})\geq g(L^{x},L^{p})$. On the LHS there is a function $f_{\gamma}(L^{x},L^{p})$ which depends on CM $\gamma$ of the tested state and Cholesky matrices $L^{x,p}$ of some witness matrices $Z_{A}^{x,p}$. The RHS of the condition, which we derive in the next section, imposes a lower bound on the LHS for all biseparable states, and it is formed by a function $g(L^{x},L^{p})$ depending only on the matrices $L^{x,p}$. The biseparability condition then can be used for testing of the presence of GME in a given CM $\gamma$ by numerical minimization of the difference $f_{\gamma}(L^{x},L^{p})-g(L^{x},L^{p})$ over the matrices $L^{x,p}$. If the minimum is negative, i.e., the biseparability condition is violated, then the state with CM $\gamma$ is inevitably GME. Here, we are interested only in a generic form of the biseparability condition capable of detecting the largest number of GME states from a minimal set of marginal CMs. Clearly, such a condition is obtained if the function $f_{\gamma}(L^{x},L^{p})$ depends on the maximum number of independent variables, i.e., if the Cholesky matrices $L^{x,p}$ corresponding to the minimal set have the maximum number of non-zero elements. Working with the positive-semidefinite matrices $Z_{A}^{x,p}$, we would have to take into account that some diagonal elements of the Cholesky matrices $L^{x,p}$ can be equal to zero [64, p. 8.3] thereby reducing the number of the variables over which we will optimize. For this reason we assume from now that the matrices $Z_{A}^{x,p}$ are positive-definite which implies that the corresponding Cholesky matrices $L^{x,p}$ possess strictly positive diagonal elements which can be expressed as [63, Theorem 4.2.5]
@@
Now, let us rewrite the first N1N-1 combinations (19) in the compact form
@@
$i=1,2,\ldots,N-1$, where the symbol $\cdot\,$” stands for the scalar product and
@@
are $(N-i)\times 1$ column vectors comprising lower triangular part without the main diagonal of the matrix $L^{\alpha}$. Hence, we see that the operators (4.1) will contain at most two non-zero terms if each of the vectors $L^{\alpha}_{i+1:N,i}$ will have at most one non-zero component. In the Appendix B we show that this will be the case if all vectors $(Z_{A}^{\alpha})_{i+1:N,i}$, $i=1,2,\ldots,N-1$ will possess at most one non-zero component and the vector (23) is the given by
$L^{\alpha}_{i+1:N,i}=\frac{(Z_{A}^{\alpha})_{i+1:N,i}}{L_{ii}^{\alpha}}.$ (24)
Making further use of the formula (16) we can write